A low-power laser used in a physics lab might have a power of 0.50mW and a beam diameter of 3.0mm. Calculate (a) the average light intensity of the laser beam, and (b) compare it to the intensity of a lightbulb producing 100-W light viewed from 2.0m. May you please show the work; thank you!
Answer #1

A) Average light Intensity is I = Power /Area = P/A = (0.5*10^-3)/(3.142*(1.5*10^-3)^2) = 70.72 W/m^2

B) I = P/A = 100/(4*3.142*2^2) = 1.989 W/m^2

So laser beam intensity is greater than light bulb